OpenStudy (anonymous):

i think i am still confused with this A and B things what about If A and B are positive constants, then find lim x-> 0 tan(Ax^2)/(Bx)^2?

OpenStudy (anonymous):

i got answer A/ B^2 any comment?

OpenStudy (solomonzelman):

Hello!

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \lim_{x\to0} \frac{\tan(Ax^2)}{Bx} }\) By applying L'Hospital's Rule I get: \(\color{#000000}{ \displaystyle \lim_{x\to0} \frac{2Ax\sec^2(Ax^2)}{B} }\) You can re-write the limit, but answer is pretty obvious, \(\color{#000000}{ \displaystyle \left(\lim_{x\to0} \frac{2A\sec^2(Ax^2)}{B} \right)\times \lim_{x\to0} x }\) So, yes you still end up with ....

OpenStudy (solomonzelman):

have you learned L'Hospital's rule?

OpenStudy (solomonzelman):

In fact, for all \(n>1\). \(\color{#000000}{ \displaystyle \lim_{x\to0} \frac{\tan(Ax^n)}{Bx} =0 }\) I can prove this by applying L'Hospital's rule: \(\color{#000000}{ \displaystyle \lim_{x\to0} \frac{nAx^{n-1}\sec^2(Ax^n)}{B} }\) And then I get, \(\color{#000000}{ \displaystyle \left( \lim_{x\to0} \frac{n\sec^2(Ax^n)}{B} \right)\times \lim_{x\to0} Ax^{n-1} =\frac{n}{B}\times 0=0 }\)

OpenStudy (solomonzelman):

If n<1, then the limit would diverge. If n=1, then following the same derivation the limit is n/B.

Directrix (directrix):

> i got answer A/ B^2 I agree. @kws942002